OpenStudy (moka414):

Find the slope of the line through the pair of points A(2,-3), P(2,9) can someone help please. thank you.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ 2}}\quad ,&{\color{blue}{ -3}})\quad B&({\color{red}{ 2}}\quad ,&{\color{blue}{ 9}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

use that same result for these too.. m=\[\frac{ (y2-y1) }{ (x2-x1) }\] here x1=x2=2,y1=-3,y2=9

OpenStudy (moka414):

I subtract both of them then?

OpenStudy (moka414):

1/0?

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (2,-3)..x1 = 2 and y1 = -3 (2,9)...x2 = 2 and y2 = 9 now sub slope(m) = (9 - (-3) / (2 - 2) slope(m) = (9 + 3) / 0 slope(m) = 12/0 When the denominator is 0, then the line is a vertical line with an undefined slope. IF the numerator would have been 0, then the line would have been horizontal with a slope of 0.

OpenStudy (moka414):

Gotcha!!! thank you.

OpenStudy (texaschic101):

sure thing :)